# Magnetic Field Line Chaos, Cantori, and Turnstiles in Toroidal Plasmas

### v1

@@ III Breakup of magnetic surfaces
An island chain forms at every rational surface that resonates with the perturbation. Realistic magnetic perturbations have many Fourier harmonics. Only in a cylinder with two symmetry directions is a single Fourier harmonic perturbation possible. (Chirikov, 1960) noted that when island chains from different rational surfaces overlap the region between the rational surfaces becomes chaotic.
IV Chaos
@@
Chaos in the sense of exponentiation allows a magnetic field $\bm{\vec{B}(\vec{x})}$ that has variations only on long spatial scales to have structures in the magnetic field lines on arbitrarily small spatial scales. As illustrated in Figure 3, the exponentiation in separation by neighboring line causes an exponentially large distortions in any tube of magnetic flux including the tubes formed by neighboring irrational magnetic surfaces. This is distortion is of fundamental importance to the physics of magnetic reconnection, Section VI, though generally ignored in the reconnection literature. Neither localized currents nor localized magnetic fields are required. Chaos is observed in curl-free magnetic fields that depend on all three spatial coordinates and are not carefully designed to avoid chaos. Nevertheless, a large exponentiation while field lines lie on perfect magnetic surfaces does not seem possible in a curl-free field.
@@
Chaotic magnetic field lines in which single field line covers the whole chaotic volume also separate exponentially when followed far enough in $\ell$. But, the existence of cantori with highly collimated turnstiles makes the implication of the first property of chaotic field lines by the second also subtle. As discussed by (MacKay, Meiss & Percival, 1984) and in the Introduction, the formation of chaotic regions in which a single line covers a volume involves cantori, which can appear to be just an irrational magnetic surface for an arbitrarily long distance $\ell$ along one of the field lines in the surface. These subtleties are also discussed in Section V and must be understood to determine the relevance of magnetohydrodynamic (MHD) simulations to actual plasma behavior.
@@
A subtlety of a field line filling a volume is illustrated by a point made by (Meiss, 2015). Although the red area in Figure 2 was formed by a single chaotic trajectory during a million iterations, that trajectory escaped from the red area after approximately a billion iterations. The implication is that the boundary of the red area was a cantorus penetrated by turnstiles that occupy only a tiny fraction of the area of the boundary. (Punjabi & Boozer, 2025) found that outermost confining magnetic surface of a stellarator has a similar subtlety. A region with a width of a few per cent of $\psi_{t}$ of the outermost confining surface confines magnetic field lines while they make transits through hundreds of thousands of stellarator periods, but they eventually strike the wall. There is presumably little difference in the behavior of the plasma on these lines than that on lines that have perfect confinement.
@@
The Brouwer fixed-point theorem of topology guarantees that field lines that close on themselves exist within any bounded chaotic field-line region. These lines can still be chaotic in having the first property of neighboring field lines with an exponential separation although obviously not the second property of covering the entire chaotic region. Although these closed lines can have neighboring lines that exponentially separate, this is not necessary because they can be the magnetic axis of a magnetic island—the line encircled by neighboring lines to form magnetic surfaces in an island. Figure 2 illustrates the existence of such islands within the red region.
@@
When an electron pressure gradient exists along chaotic magnetic field lines, (Boozer, 2024b) showed that the electric field required for quasi-neutrality gives $\vec{E}\times\vec{B}$ flows that lead to plasma transport at a Bohm-like rate. A variation in the electric potential, $\partial\Phi/\partial\ell=e\partial P_{e}/\partial\ell$, is required to keep the rapidly moving electrons together with the slowly moving ions. The exponentially large dependence of the separation of neighboring field lines with $\ell$ leads to increasing $\vec{E}\times\vec{B}$ flows until a balance in electron and ion transport is achieved. The required Bohm-like level of transport across the magnetic field lines is important for sweeping in impurities in tokamak disruptions and for transport in non-resonant stellarator divertors.
@@
Velocities can also be chaotic by having chaotic stream lines. In a chaotic flow, each streamline, $\bm{\vec{x}(t)}$ with $\bm{d\vec{x}/dt=\vec{u}(\vec{x},t)}$, has neighboring streamlines that increase their separation exponentially with time. A chaotic magnetic field requires a dependence on all three spatial coordinates. A chaotic magnetic field has chaotic magnetic field lines, which are defined at fixed instants in time. However, a chaotic flow requires a dependence on only two spatial coordinates when the flow is time dependent. The failure to appreciate the implications explains the origins of fundamental misconceptions in the theory of magnetic reconnection, Section VI. As discussed in (Boozer, 2025b) when the flow velocity of magnetic field lines is chaotic, the ideal evolution of a magnetic field in a two-spatial coordinate system gives an exponential increase of the field strength. When the field depends on all three spatial coordinates, a chaotic flow velocity makes the field lines chaotic while producing only a moderate increase in the magnetic field strength.
V Cantori and turnstiles
@@
In an evolving $\vec{B}$, the last magnetic surface to break is the surface with iota most difficult to approximate by a rational. As an irrational surface breaks, it forms a cantorus, which is a misnomer since it need not have the self-similarity at all spatial scales of cantor sets. Cantori can also exist in static magnetic fields in the region in which confining magnetic surfaces transition into open magnetic field lines that strike the walls, as in a non-resonant stellarator divertor.
@@
Conventionally subtleties of magnetic field line evolution are understood using Poincaré plots. (Boozer, 2025a) showed far more information can be obtained by Fourier decomposing $R(\varphi)$ and $Z(\varphi)$ of magnetic field lines that are followed in ordinary $(R,\varphi,Z)$ cylindrical coordinates to find surfaces on which magnetic field lines lie or near which they linger. Knowing these surfaces, $\vec{B}\cdot\hat{n}$ can be determined, cantori identified, the magnitude and collumation of the magnetic flux that leaks through these surfaces in turnstiles can found.
On a magnetic surface, the magnetic poloidal angle obeys $\Theta=\iota_{s}\varphi$, where $\iota_{s}$ is a constant, the rotational transform. The expansion of $Z(\Theta,\varphi)=\sum_{mn}Z_{mn}\cos(m\Theta-n\varphi)$ can be found using a Gaussian function of width $w$, which satisfies $\int_{0}^{\infty}G(\varphi)\cos(\Omega\varphi)d\varphi=\exp(-(w\Omega)^{2}/2)$. The function $\bm{Z_{f}(\omega)\equiv 2\int_{0}^{\infty}G(\varphi)\cos(\omega\zeta)Z(\varphi)d\varphi}$ can be found by a Fast Fourier Transform. When $w|\omega_{mn}|>>1$, where $\omega_{mn}\equiv m\iota_{s}-n$, the function $Z_{f}(\omega)$ has the form $\sum_{mn}Z_{mn}\exp(-(w^{2}/2)(\omega-\omega_{mn})^{2})$, which determines the Fourier coefficients and $\iota_{s}$. For simplicity, it has been assumed only sine or cosine terms are required for the Fourier decomposition for $R$ and $Z$.
@@
The inverse of the Jacobian, $\vec{\nabla}\psi_{t}\times\vec{\nabla}\Theta)\cdot\vec{\nabla}\varphi=2\pi\vec{B}\cdot\vec{\nabla}\varphi$, and the toroidal field $B_{t}=R\vec{B}\cdot\vec{\nabla}\varphi$ is almost constant on a surface.
@@
Reconnection with neighboring surfaces becomes logarithmically easier the larger the ratio of the maximum to the minimum separation on the surface. The larger $\sigma_{s}$ is the more localized is reconnection to points on the surface where the minimum separation $\Delta_{min}$ is located.
One subtlety is that there is no guarantee that surfaces do not overlap. When there is no chaos, so the widths of the Gaussian lines decrease as $w$ increases, ensuring surfaces do not overlap should be a matter of numerical resolution. Nevertheless, fast disruptions in high temperature tokamaks could in principle involve ratios of $\Delta_{max}/\Delta_{min}\sim 10^{20}$, which seems daunting. Of course far smaller ratios should clearly identify the points on the surface at which reconnection takes place and turnstiles form. The actual locations of turnstiles are the places at which $\vec{B}\cdot\hat{n}\neq 0$, where the unit normal to the surface is
@@
The turnstiles that form the divertor legs of non-resonant divertors strike the walls in stripes in remarkably robust locations (Bader et al., 2018). (Boozer & Punjabi, 2018) and (Garcia et al., 2025) found the stripes have a width of order a percent on the wall area.
VI Magnetic reconnection
Three concepts are needed to understand magnetic reconnection: (1) magnetic field line chaos, (2) the magnetic field line velocity, and (3) and its chaos. The magnetic field line velocity $\vec{u}_{\bot}$ is distinct from the plasma velocity $\vec{v}$.
VI.1 Magnetic field line velocity $\vec{u}_{\bot}$
@@
The loop voltage in a region of spatially bounded field lines is defined by Equation (3). When $\vec{\nabla}V_{\ell}=0$ the magnetic evolution is ideal. Faraday’s law then implies the evolution of Clebsch coordinates $\alpha$ and $\beta$ where $\vec{B}=\vec{\nabla}\alpha\times\vec{\nabla}\beta$ is consistent with $d\alpha/dt=\partial\alpha/\partial t+\vec{u}_{\bot}\cdot\vec{\nabla}\alpha=0$ and $d\beta/dt=\partial\beta/\partial t+\vec{u}_{\bot}\cdot\vec{\nabla}\beta=0$. Since $\alpha$ and $\beta$ define a magnetic field line, $\vec{u}_{\bot}$ is the velocity of the magnetic field lines when the gradient of the loop voltage is zero. This is in essence the proof given in (Newcomb, 1958). The physical interpretation of $\vec{u}_{\bot}$ is subtle when $\vec{\nabla}V_{\ell}\neq 0$, but $\vec{u}_{\bot}$ is still defined and for simplicity of language can be called the magnetic field line velocity. The fact that $\vec{u}_{\bot}$ depends on the frame of reference that is used, which can change if canonical transformations are made to $(\psi_{t},\theta,\varphi)$ as the system evolves, has created confusion about the existence of $\vec{u}_{\bot}$.
Equation (17) for $\vec{E}$ follows from mathematics alone—the representation of vectors in three-space. Although (Boozer, 2025b) has shown that Equation (17) is far more generally valid, the proof is more obvious when $\vec{B}\cdot\vec{\nabla}\varphi\neq 0$. An arbitrary vector $\vec{E}$ is three-space can be represented in terms of another arbitrary vector $\vec{B}$ when all three components of $\vec{E}$ can be represented. The component $\vec{B}\cdot\vec{E}=-\vec{B}\cdot\vec{\nabla}\Phi+V_{\ell}\vec{B}\cdot\vec{\nabla}\varphi$ is equivalent to
@@
Locally this equation for $\Phi$ along a magnetic field line always has a solution. However, a loop voltage $V_{\ell}$ is generally required to obtain periodicity in $\varphi$ in a torus or when there are two boundary conditions on $\Phi$ as there generally are when a field lines strike walls. The velocity $\vec{u}_{\bot}$ can always be chosen so the two components of $\vec{E}$ perpendicular to $\vec{B}$ are represented.
@@
(Aref, 1984) illustrated the exponentially large effect on mixing in fluids when the velocity with which they are stirred is chaotic. (Tang & Boozer, 1999) used Lagrangian coordinates to show the effective spatial diffusion increases exponentially in time.
VI.2 Magnetic reconnection when $\vec{B}$ is nearly ideal
When $\vec{B}$ is ideal and $\vec{u}_{\bot}$ is chaotic, the surfaces that enclose a volume of magnetic flux increase their surface area exponentially without limit even though the enclosed magnetic flux is a constant, Figure 3. (Boozer and Elder 2021) gave an example of smooth large-scale non-turbulent magnetic field line flow $\vec{u}_{\bot}$ that is chaotic. Their example is particularly interesting because it puts no helicity into the system. As they discuss, unlike magnetic energy, the dissipation of magnetic helicity cannot be significantly enhanced by turbulence. In a system like the solar corona in which magnetic flux tubes are driven by plasma motion in the photosphere, helicity buildup cannot be bounded other than by the ejection of the loop from the sum, a coronal mass ejection.

### v2

@@ 3 Breakup of magnetic surfaces
An island chain forms at every rational surface that resonates with the perturbation. Realistic magnetic perturbations have many Fourier harmonics. Only in a cylinder with two symmetry directions is a single Fourier harmonic perturbation possible. (Chirikov:1960) noted that when island chains from different rational surfaces overlap the region between the rational surfaces becomes chaotic.
4 Chaos
@@
Chaos in the sense of exponentiation allows a magnetic field $\vec{B}(\vec{x})$ that has variations only on long spatial scales to have structures in the magnetic field lines on arbitrarily small spatial scales. As illustrated in Figure 3, the exponentiation in separation of neighboring lines causes an exponentially large distortions in any tube of magnetic flux including the tubes formed by neighboring irrational magnetic surfaces. This distortion is of fundamental importance to the physics of magnetic reconnection, Section 6, though generally ignored in the reconnection literature. Neither localized currents nor localized magnetic fields are required. Chaos is observed in curl-free magnetic fields that depend on all three spatial coordinates and are not carefully designed to avoid chaos. Nevertheless, a large exponentiation while field lines lie on perfect magnetic surfaces does not seem possible in a curl-free field.
@@
Chaotic magnetic field lines in which a single field line covers the whole chaotic volume also separate exponentially when followed far enough in $\ell$. But, the existence of cantori with highly collimated turnstiles makes the implication of the first property of chaotic field lines by the second also subtle. As discussed by (MacKay:1984) and in the Introduction, the formation of chaotic regions in which a single line covers a volume involves cantori, which can appear to be just an irrational magnetic surface for an arbitrarily long distance $\ell$ along one of the field lines in the surface. These subtleties are also discussed in Section 5 and must be understood to determine the relevance of magnetohydrodynamic (MHD) simulations to actual plasma behavior.
@@
A subtlety of a field line filling a volume is illustrated by a point made by (Meiss:2015). Although the red area in Figure 2 was formed by a single chaotic trajectory during a million iterations, that trajectory escaped from the red area after approximately a billion iterations. The implication is that the boundary of the red area was a cantorus penetrated by turnstiles that occupy only a tiny fraction of the area of the boundary. (Punjabi:2025) found that outermost confining magnetic surface of a stellarator has a similar subtlety. A region with a width of a few per cent of the $\psi_{t}$ enclosed by the outermost confining surface confines magnetic field lines while they make transits through hundreds of thousands of stellarator periods, but they eventually strike the wall. There is presumably little difference in the behavior of the plasma on these lines than that on lines that have perfect confinement.
@@
The Brouwer fixed-point theorem of topology guarantees that field lines that close on themselves exist within any bounded chaotic field-line region. These lines can still be chaotic in having the first property of neighboring field lines with an exponential separation although obviously not the second property of covering the entire chaotic region. Although these closed lines can have neighboring lines that exponentially separate, this is not necessary because they can be the magnetic axis of a magnetic island. Figure 2 illustrates the existence of such islands within the red region.
@@
When the electron pressure, $P_{e}$, has a gradient along chaotic magnetic field lines, (Boozer:2024:Electric) showed that the electric field required for quasi-neutrality gives $\vec{E}\times\vec{B}$ flows that lead to plasma transport at a Bohm-like rate. A variation in the electric potential, $en\partial\Phi/\partial\ell=\partial P_{e}/\partial\ell$, where $n$ is the electron number density, is required to keep the rapidly moving electrons together with the slowly moving ions. The exponentially large dependence of the separation of neighboring field lines with $\ell$ leads to increasing $\vec{E}\times\vec{B}$ flows until a balance in electron and ion transport is achieved. The required Bohm-like level of transport across the magnetic field lines is important for sweeping in impurities in tokamak disruptions and for transport in non-resonant stellarator divertors.
@@
Velocities can also be chaotic by having chaotic streamlines. In a chaotic flow, each streamline, $\vec{x}(t)$ with $d\vec{x}/dt=\vec{u}(\vec{x},t)$, has neighboring streamlines that increase their separation exponentially with time. A chaotic magnetic field requires a dependence on all three spatial coordinates. A chaotic magnetic field has chaotic magnetic field lines, which are defined at fixed instants in time. However, a chaotic flow requires a dependence on only two spatial coordinates when the flow is time dependent. The failure to appreciate the implications explains the origins of fundamental misconceptions in the theory of magnetic reconnection, Section 6. As discussed in (Boozer:2025:Magnetic), when the flow velocity of magnetic field lines is chaotic, the ideal evolution of a magnetic field in a system with only two spatial coordinates produces an exponential increase of the field strength. When the field depends on all three spatial coordinates, a chaotic flow velocity makes the field lines chaotic while producing only a moderate increase in the magnetic field strength.
5 Cantori and turnstiles
@@
In an evolving $\vec{B}$, the last magnetic surface to break is the surface with iota most difficult to approximate by a rational. As an irrational surface breaks, it forms a cantorus. Cantori can also exist in static magnetic fields in the region in which confining magnetic surfaces transition into open magnetic field lines that strike the walls, as in a non-resonant stellarator divertor.
@@
5.1 Efficient way to obtain information about cantori and turnstiles
Conventionally subtleties of magnetic field line evolution are understood using Poincaré plots. (Boozer:2025:Efficient) has shown that far more information can be obtained by Fourier decomposing $R(\varphi)$ and $Z(\varphi)$ of magnetic field lines that are followed in ordinary $(R,\varphi,Z)$ cylindrical coordinates to find surfaces on which magnetic field lines lie or near which they linger. Knowing these surfaces, $\vec{B}\cdot\hat{n}$ can be determined, cantori identified, the magnitude and collumation of the magnetic flux that leaks through these surfaces in turnstiles can be found.
On a magnetic surface, the magnetic poloidal angle obeys $\Theta=\iota_{s}\varphi$, where $\iota_{s}$ is a constant, the rotational transform. The expansion of $Z(\Theta,\varphi)=\sum_{mn}Z_{mn}\cos(m\Theta-n\varphi)$ can be found using a Gaussian function of width $w$, which satisfies $\int_{0}^{\infty}G(\varphi)\cos(\Omega\varphi)d\varphi=\exp(-(w\Omega)^{2}/2)$. The function $Z_{f}(\omega)\equiv 2\int_{0}^{\infty}G(\varphi)\cos(\omega\zeta)Z(\varphi)d\varphi$ can be found by a Fast Fourier Transform. When $w|\omega_{mn}|>>1$, where $\omega_{mn}\equiv m\iota_{s}-n$, the function $Z_{f}(\omega)$ has the form $\sum_{mn}Z_{mn}\exp(-(w^{2}/2)(\omega-\omega_{mn})^{2})$, which determines the Fourier coefficients and $\iota_{s}$. For simplicity, it has been assumed only sine or cosine terms are required for the Fourier decomposition for $R$ and $Z$.
@@ 5.1 Efficient way to obtain information about cantori and turnstiles
The inverse of the Jacobian, $(\vec{\nabla}\psi_{t}\times\vec{\nabla}\Theta)\cdot\vec{\nabla}\varphi=2\pi\vec{B}\cdot\vec{\nabla}\varphi$, and the toroidal field $B_{t}=R\vec{B}\cdot\vec{\nabla}\varphi$ is almost constant on a surface.
@@
Reconnection with neighboring surfaces becomes exponentially easier the larger the ratio of the maximum to the minimum separation on the surface. The larger $\sigma_{s}$ is the more localized is reconnection to points on the surface where the minimum separation $\Delta_{min}$ is located.
One subtlety is that there is no guarantee that surfaces do not overlap. When there is no chaos, so the widths of the Gaussian lines decrease as $w$ increases, ensuring surfaces do not overlap should be a matter of numerical resolution. Nevertheless, fast disruptions in high temperature tokamaks could in principle involve ratios of $\Delta_{max}/\Delta_{min}\sim 10^{10}$, which seems daunting. Of course far smaller ratios should clearly identify the points on the surface at which reconnection takes place and turnstiles form. The actual locations of turnstiles are the places at which $\vec{B}\cdot\hat{n}\neq 0$, where the unit normal to the surface is
@@
5.2 Subtleties of cantori and turnstiles in non-resonanant divertors
An intuitive understanding of subtleties of cantori and turnstiles can be obtained from a study of the localization of the plasma flows in a stellarator non-resonant divertor. This can be done by following field lines from the point where they enter the plasma chamber until they leave. The strong helical fields outside of the magnetic surfaces of a stellarator cause almost all magnetic field lines to exit the plasma chamber within approximately a stellarator period of where they entered. However, there are narrow helical stripes on the walls, which take up only a few percent of the wall area, with entry points that give lines that make many toroidal transits before restriking the wall. Indeed, there are entry points within these stripes from which field lines make an arbitrarily large number of toroidal transits (Punjabi:2025). The field lines that come closest to the confined plasma take the largest number of toroidal transits between wall entry and departure. These lines form a tube of magnetic flux that can be filled with plasma by cross-field diffusion and produce the divertor. This flux tube is roughly circular but strikes the wall in a helical stripe since the magnetic field component tangential to the stellarator walls is strong in comparison to the normal field.
The behavior of field lines in non-resonant divertors is understandable due the magnetic field just outside the last confining magnetic having a number of important cantori with the cantori closer to the outermost confining magnetic surface having narrower turnstiles. Once a field line passes through the outermost important turnstile, which occupies only a few percent of the cantorus area, it takes many toroidal transits to find an exit hole which must have the same small magnetic flux as the entry turnnstile. Once the field line exits that turnstile, the line strikes the wall in a fraction of a stellarator period, which is too short a distance for the exponentiation of the chaos to destroy the collimation produced by the small hole defined by the turnstile.
While the field line is making toroidal transits between the entry and the exit turnstile of the outermost important cantorus, it can intercept an even smaller turnstile of a cantoris that is closer to the last confining magnetic surface. These lines become trapped for an even larger number of toroidal transits before they intercept an existing turnstile. The strong correlation between the closeness with which a field line approaches the outermost confining magnetic surface and the number of toroidal transits the line makes is illustrated in (Garcia:2025) and (Punjabi:2025).
The location of the outermost important turnstile is determined by a strong curvature in the outermost magnetic surface, which are called helical edges. They resemble the X-lines of a tokamak divertor. Unlike X-lines, which are toroidally continuous in axisymmetry and require the rotational transform to be zero, helical edges extend only over a fraction of stellarator period. A sharp helical edge on a magnetic surface, or an X-line of a tokamak divertor, can only be produced by currents for more distant than the radius of curvature of the magnetic surface when the edge is parallel to a magnetic field line. Since the helical edge is not a closed curve, it does not define the rotational transform of the surface. (Bader:2018) has shown the locations on the wall that are struck by the field lines of a non-resonant divertor are robust against plasma changes and are determined by strong variation in the curvature of the outermost magnetic surface.
6 Magnetic reconnection
Three concepts are needed to understand magnetic reconnection: (1) magnetic field line chaos, (2) the magnetic field line velocity, $\vec{u}_{\bot}$, and (3) and its chaos. The magnetic field line velocity $\vec{u}_{\bot}$ is distinct from the plasma velocity $\vec{v}$.
6.1 Magnetic field line velocity $\vec{u}_{\bot}$
@@
The loop voltage in a region of spatially bounded field lines is defined by Equation (3). When $\vec{\nabla}V_{\ell}=0$ the magnetic evolution is ideal. Faraday’s law then implies the evolution of Clebsch coordinates $\alpha$ and $\beta$ where $\vec{B}=\vec{\nabla}\alpha\times\vec{\nabla}\beta$ is consistent with $d\alpha/dt=\partial\alpha/\partial t+\vec{u}_{\bot}\cdot\vec{\nabla}\alpha=0$ and $d\beta/dt=\partial\beta/\partial t+\vec{u}_{\bot}\cdot\vec{\nabla}\beta=0$. Since $\alpha$ and $\beta$ define a magnetic field line, $\vec{u}_{\bot}$ is the velocity of the magnetic field lines when the gradient of the loop voltage is zero. This is in essence the proof given in (Newcomb:1958). The physical interpretation of $\vec{u}_{\bot}$ is subtle when $\vec{\nabla}V_{\ell}\neq 0$, but $\vec{u}_{\bot}$ is still defined and for simplicity of language can be called the magnetic field line velocity. The fact that $\vec{u}_{\bot}$ depends on the frame of reference that is used, which can change if canonical transformations are made to $(\psi_{t},\theta,\varphi)$ as the system evolves, has created confusion about the existence of $\vec{u}_{\bot}$.
Equation (17) for $\vec{E}$ follows from mathematics alone—the representation of vectors in a three dimensional space. Although (Boozer:2025:Magnetic) has shown that Equation (17) is far more generally valid, the proof is more obvious when $\vec{B}\cdot\vec{\nabla}\varphi\neq 0$. An arbitrary vector $\vec{E}$ in three-space can be represented in terms of another arbitrary vector $\vec{B}$ when all three components of $\vec{E}$ can be represented. The component $\vec{B}\cdot\vec{E}=-\vec{B}\cdot\vec{\nabla}\Phi+V_{\ell}\vec{B}\cdot\vec{\nabla}\varphi$ is equivalent to
@@
Locally this equation for $\Phi$ along a magnetic field line always has a solution with $V_{\ell}=0$. However, a loop voltage $V_{\ell}$ is generally required to make $\Phi$ a periodic function $\varphi$ in a torus or when there are two boundary conditions on $\Phi$ to be satisfied as there generally are when field lines strike walls. The velocity $\vec{u}_{\bot}$ can always be chosen so the two components of $\vec{E}$ perpendicular to $\vec{B}$ are represented.
@@
(Aref:1984) illustrated the exponentially large effect on mixing in fluids when the velocity with which they are stirred is chaotic. (Tang:1999) used Lagrangian coordinates to show the effective spatial diffusion increases exponentially in time.
6.2 Magnetic reconnection when $\vec{B}$ is nearly ideal
When $\vec{B}$ is ideal and $\vec{u}_{\bot}$ is chaotic, the surfaces that enclose a volume of magnetic flux increase their surface area exponentially without limit even though the enclosed magnetic flux is a constant, Figure 3. (Boozer and Elder 2021) gave an example of smooth large-scale non-turbulent magnetic field line flow $\vec{u}_{\bot}$ that is chaotic. Their example is particularly interesting because it puts no helicity into the system. As they discuss, unlike magnetic energy, the dissipation of magnetic helicity cannot be significantly enhanced by turbulence. In a system like the solar corona in which magnetic flux tubes are driven by plasma motion in the photosphere, helicity buildup cannot be bounded other than by the ejection of the loop from the sun, a coronal mass ejection.